OpenStudy (anonymous):

integral calculus: determine the length of the arc of the curve y=e^x from x=0 to x=1?

OpenStudy (anonymous):

L = ∫ sqrt(1 + (dy/dx)^2) dx

OpenStudy (anonymous):

i dont get it maam

OpenStudy (freckles):

do you know what dy/dx means ?

OpenStudy (anonymous):

yes its the derivative

OpenStudy (freckles):

ok well as you can see the formula needs you to calculate that

geerky42 (geerky42):

To find length of arc, you use this: \[\int\sqrt{1+\left(\dfrac{\mathrm dy}{\mathrm dx}\right)^2}\mathrm dx\] Here, we have \(y = e^x\) Obviously, \(\dfrac{\mathrm dy}{\mathrm dx} = e^x\) So you have to evaluate \[\int\sqrt{1+\left(e^x\right)^2}\mathrm dx = \int\sqrt{1+e^{2x}}\mathrm dx \]

OpenStudy (anonymous):

thanks sir

OpenStudy (anonymous):

thanks for the clearer explanation @geerky42